You invested $8000 in two accounts paying 3% and 8% annual interest, respectively. If the total interest earned for the year was $490, how much was invested at each rate?
Answer by mananth(16949)

let the amount invested at 3% be x
the amount invested at 8% will be 8000-x
0.03x+0.08(8000-x)=490
0.03x + 640 - 0.08x=490
-0.05x= -150
x= -150 / -0.05
x=3000 which is amount invested at 3%
Balance $5000 at *%